Consider a compact disc on which the pits used to store the digital information measure 850 nm long by 500 nm wide. Visible light has the following wavelengths: 380-420 nm for violet; 420-440 nm for indigo; 440 nm-500 nm for blue; 500-520 nm for cyan; 520-565 nm for green; 565-590 nm for yellow; 590-625 nm for orange, 625-740 nm for red. Which of these colors, if any, could be used for building the pits on the disc? List any that will work.
Answer #1

Solution :-

Given data,

Wavelength ,

Voilet = 380- 420nm

Indigo = 420-440nm

Blue = 440-500 nm

Cyan = 500-520 nm

Green = 520-565 nm

Yellow = 565-590 nm

Orange = 590-625 nm

Red. = 625-740 nm

For storing information the waves should form a standing wave on the pits.

Hence the wavelength should be integral multiples of the dimensions.

Hence

Indigo i.e. 420-440 nm will form at 850 nm side with n =2

And Blue and Cyan will form at 500nm